40m QIL Cryo_Lab CTN SUS_Lab TCS_Lab OMC_Lab CRIME_Lab FEA ENG_Labs OptContFac Mariner WBEEShop
  40m Log, Page 234 of 348  Not logged in ELOG logo
ID Date Author Type Category Subject
  5757   Fri Oct 28 15:33:06 2011 JenneUpdateComputersNifty screen generator

Suresh showed me a cool script that Mirko made, but didn't elog about.

You tell the script what filter banks you want, and it creates a screen for each with a bunch of different filter module display formats.  Then you can copy the format you like into the actual screen you're modifying. 

Currently PEM, LSC and IOO (and maybe others?) have "fmX" folders inside their medm/c1.../master folders.  For each subsystem, we need to copy this folder, and modify the generic .adl file so that it puts in the correct subsystem letters.  Once this is done, you can just run the generateFMscreens.py after putting in your filter bank names.

  5756   Fri Oct 28 14:56:02 2011 JenneUpdateCDSCSS/BOY installed on pianosa

Quote:

I've installed Control System Studio (CSS) on pianosa, from the version 3.0.2 Red Hat binary zip.  It should be available as "css" from the command line.

CSS is a new MEDM replacement. It's output is .opi files, instead of .adl files.  It's supposed to include some sort of converter, but I didn't play with it enough to figure it out.

Please play around with it and let me know if there are any issues.

links:

 So far I've only given it about half an hour of my time, but it is *really* frustrating so far.  There don't seem to be any instructions on how to tell it what our channels are / how to link CSS to our EPICS databases.  Or, the instructions that are there say "do it!", but they neglect to mention how...  Also, there exists (maybe?) an ADL->BOY converter, but I can't find any buttons to click, or how to import an .adl, or what I'm supposed to do.  Also, it's not clear how to get to the editor to start making screens from scratch. 

It looks like it has lots of nifty indicators and buttons, but I would have felt better if I had been able to do anything.

Another thing that is going to be a problem:  the Shell Command button that we use all over the place in our MEDM screens is not supported by this program.  It's listed in the "limitations" of the ADL2BOY converter.  This may kill the CSS program immediately.  Jamie: did Rolf/anyone mention a game plan for this?  It's super nice to be able to run scripts from the screens.

Moral of the story:  I'm annoyed, and going to continue making my OAF screens in MEDM for now.

  5755   Fri Oct 28 12:47:38 2011 jamieUpdateCDSCSS/BOY installed on pianosa

I've installed Control System Studio (CSS) on pianosa, from the version 3.0.2 Red Hat binary zip.  It should be available as "css" from the command line.

CSS is a new MEDM replacement. It's output is .opi files, instead of .adl files.  It's supposed to include some sort of converter, but I didn't play with it enough to figure it out.

Please play around with it and let me know if there are any issues.

links:

  5754   Fri Oct 28 05:17:13 2011 kiwamuUpdateLSClocking activity : PZT1 is still railing

Status update on the LSC activity:

 To see how good/bad the beam pointing is, I locked the Y arm with POY11.
Then I ran the ASS servo to automatically correct the alignment of the ITMY and ETMY suspensions and also the beam pointing.
The result is that the PZT1_X is still railing to the negative side.
Due to it the transmitted light from the Y arm is about 0.6 or so which is supposed to be 1 if the beam pointing is perfect.
The EPICS value of PZT1_X is at the minimum of -10 and the ASS servo tried to push it more negative side.
 
 Tomorrow night I will intentionally introduce offsets in the MC suspensions to avoid the railing.
The goal will be a scan of the incident beam while measuring the recycling gain.
  5753   Fri Oct 28 04:57:00 2011 kiwamuUpdateLSCPOX11 demod board broken
The POX11 demodulation board is broken. It needs to be fixed in the daytime tomorrow.
It only outputs the Q signal and nothing is coming out from the I output.


(Some stuff checked)
 [OK] ADCs
 [OK] Whitening filters looked fine. Their gains were controllable from EPICS.
 [OK] Connection between the POX11 RFPD and demodulator box
 [OK] The Q signal showed the PDH signal of the X arm with an amplitude of about 200 counts, which almost is the same as that of POY11.
 [NOT GOOD] The I signal from the demod board
 [OK] Outputs cables, which send the I and Q signals from the demod board to WFs are fine.
         
(Test on the POX11 demod board)
 As usual, a test signal whose frequency is shifted by a little bit from that of LO was injected to the RF input of the board to see if the circuit is working.
The I signal didn't show up and there were no signals even in the monitor LEMO output.
Something is wrong in the I signal demodulation path on the circuit board.

Here is an actual time series of the I and Q signals in dataviewer. The I signal outputs just junk while the Q showed a nice sine curve.
Screen_shot_2011-10-28_at_2.26.02.png

  5752   Fri Oct 28 03:42:50 2011 kiwamuUpdateLSCITMX table needs to be refined

(POX)

The POX beam had been 80% clipped at a black glass beam dump of the POX11 RFPD.

I steered the first mirror in the POX path to fix the clipping. Then the beam was realigned onto the RFPD.

However the beam is still very close to the black glass, because the incident angle to the second mirror is not 45 deg .

We need to refine the arrangement of the POX11 optics a bit more so that the beam will never be clipped at the black glass.

 

(POP)

 The POP optics also need to be rearranged to accommodate one more RFPD.

Additionally Rana, Suresh and I discussed the possible solutions of POP22/110 and decided to install a usual PD (PDA10A or similar) instead of a custom-made.

So a plan for the POP detectors will be something like this:

        + design an optical layout.

        + buy a 2 inch lens whose focul length is long enough (#5743)

        + rearrange the optics and install POP22/110

        + lay down a long SMA cable which sends the RF signal from POP22/110 to the LSC rack.

        + install a power splitter just before the demod board so that the signal is split into the 22MHz demoad board and 110MHz demod board.

           => make sure we have a right splitter for it.

        + install a band pass filter after the power splitter in each path.

           => A 22MHz band pass filter is already in hand. Do we have 110MHz band pass filter somewhere in the lab ?

The picture here shows the latest configuration on the ITMX table.

ITMXtable.png

Quote from #5743

RF photo diodes POP55 and POX11 are installed. The beams are aligned to the photo diodes. 

  5751   Fri Oct 28 03:12:37 2011 SureshUpdateIOOMC2 realigned to align MC to PSL

Around 6PM on the 27th, I found that the C1:IOO-MC_RFPD_DCMON had risen to about 2.5V.   I checked the trend of MC2 sensors and found that  between 2PM and 6PM, MC2 had drifted in a strange way.  And also that the alignment had grown worse over several days.   I also noticed that the spot on the MC2F camera had shifted to the left.

I attempted to correct the alignment (decrease the C1:IOO-MC_RFPD_DCMON to ~0.5V ) by just moving the MC2 and succeeded!! So it is quite likely that most of the slow MC drift is arising due to MC2 table drift.

MC2_Drift_20111027.png

 

I decided to try and close the MC2_TRANS QPD to MC2 loops separately to see if MC alignment becomes stable  But several screens needed to be fixed before we could try anything.   So I fixed C1IOO_WFS_MASTER,  C1IOO_WFS_INMATRIX and ...OUTMATRIX screens.  Deleted the older ones to avoid confusion.

In this process I noticed that the directory of $screens$/c1mcs/master contains copies of older C1SUS_MC1 , 2. and 3 screens which look very similar to the new autogenerated screens.  Some of the links in the WFS screens were pointing to the old screens.  I have redirected the links and I will delete these in a couple of days, if no one objects.

 

 

 

  5750   Fri Oct 28 02:41:18 2011 SureshMetaphysicselogelog unresponsive: restarted

Elog did not respond despite running the /cvs/cds/caltech/elog/start-elog.csh  script two times.  

It worked the after the third restart. 
 

  5749   Fri Oct 28 01:13:17 2011 kiwamuUpdateSUSITMX oplev : iris fully opened

I found that the sum of the ITMX oplev signals had gone down to zero yesterday.

I checked the ITMX table and found two iris on the He-Ne laser path were blocking the beam on their apertures.

I guess this is because we were working around there for installation of POP/POX and may have touched some of the oplev optics.

Then I fully opened the apertures of those two iris and the sum went back to nominal of 600 counts.

  5748   Fri Oct 28 00:53:39 2011 ranaUpdateElectronicsPOP 22/110 Design

The attached PDF shows a possible gain / input noise config for the POP 22/110 that we would use to detect the RF power in the DRMI. Design is in the SVN.

If Kiwamu/Jenne say that this has good enough sensing noise for the lock triggering than we will build it. This is using a 2mm diode.

If we can get away with 1 mm, we might as well use a PDA10CF for now.

Attachment 1: poy22110.pdf
poy22110.pdf poy22110.pdf
  5747   Thu Oct 27 18:00:38 2011 kiwamuSummaryLSCOffsets in LSC signals due to the RFAMs : Optickle simulation

The amount of offsets in the LSC signals due to the RFAMs have been estimated by an Optickle simulation.

The next step is to think about what kind of effects we get from the RFAMs and estimate how much they will degrade the performance.

(Motivation)

  We have been having relatively big RFAM sidebands (#5616), which generally introduce unwanted offsets in any of the LSC demodulated signals.
The motivation was that we wanted to estimate how much offsets we've been having due to the RFAMs.
The extreme goal is to answer the question : 'How big RFAMs do we allow for operation of the interferometer?'.
Depending on the answer we may need to actively control the RFAMs as already planed (#5686).
Since the response of the interferometer is too complicated for analytic works, so a numerical simulation is used.
 

(Results : Offsets in LSC error signals)

PRCL_200.png

 

MICH_200.png

 SRCL_200.png

  Figure: Offsets in unit of meter in all the LSC demodulated signals.  Y-axis is the amount of the offsets and the X-axis represents each signal port.
In each signal port, the signals are classified by color.
(1) Offsets in the PRCL signal. (2) Offsets in the MICH signal. (3) Offsets in the SRCL signal.
 
 
Roughly the signals showed offsets at a 0.1 nm level.
The numerical error was found to be about 10-10 nm by running the same simulation without the AM sidebands.
Here is a summary of the amount of the offsets:
 
    offsets [nm] (1f signal port)  offsets [nm] (3f signal port)  biggest offsets [nm] (signal port)
PRCL       0.3 (REFL11)       0.2 (REFL33)     1 (REFL55)
MICH      0.00009 (AS55)       0.8 (REFL33)     7 (POP11)
SRCL      0.1 (REFL55)       0.1 (REFL165)     40 (POX11)
In the SRCL simulation  REFL11I, REFL11Q, POP11I, POP11Q and POX11I didn't show any zero crossing points within 100 nm range around the resonance.
It is because that the SRCL doesn't do anything for the 11MHz sidebands. So it is the right behavior.
However POX11 was somewhat sensitive to the SRCL motion and showed a funny signal with a big offset.
 

(Simulation setup)

I applied the current PM/AM ratio according to the measurement (#5616, #5519)
The modulation indices used in the simulation are :
    + PM index in 11MHz = 0.17
    + PM index in 55MHz = 0.14
    + AM index in 11MHz = 0.17 / 200 = 8.5x10-4
    + AM index in 55MHz = 0.14 / 200 = 7.0x10-4
Note that the phases of the AM and PM sidebands are the same.

For clarity, I also note the definition of PM/AM ratio as well as how the first order upper sideband looks like.

ratio.png

upper.png
 

The optical parameters are all at ideal length although we may want to check the results with more realistic parameters:
    + No arm cavities
    + PRCL length = 6.75380
    + SRCL length = 5.39915
    + Schnupp asymmetry = 3.42 cm
    + loss in each optic = 50 ppm
    + PRCL = resonant for 11 and 55MHz
    + MICH = dark fringe
    + SRCL = resonant for 55 MHz
The matlab script will be uploaded to the cvs server.

Quote from #5686
  8. In parallel to those actions, figure out how much offsets each LSC error signal will have due to the current amount of the RFAMs.
    => Optickle simulations.

  5746   Thu Oct 27 16:09:37 2011 kiwamuUpdateLSCRF distribution box : two more 55MHz available

The diagram of the RF distribution box has been updated according to the modification ( #5744).

Both pdf and graffle files are available on the 40m svn : https://nodus.ligo.caltech.edu:30889/svn/trunk/suresh/40m_RF_upgrade/

Here shows the latest version of the diagram.

RF_Distribution_Box.png

Quote from #5744

I will update the diagram of the RF distribution box (#4342) tomorrow.

  5745   Thu Oct 27 03:32:45 2011 KojiSummaryIOORFAM monitor progress

[Suresh, Mirko, Koji]

A cable from the stochmon box to the cross connect for the EPICS ADCs is installed.

The power supply and the signal outputs are concentrated in a single DSub 9pin connector
that is newly attached on the box.

The connection from the stochmon side of the cable and the EPICS value was confirmed.
The calibration of them looks fine.

To do:
- Once the stochmon box is completed we can immediately test it.
- The EPICS channel names are still as they were. We need to update the database file of c1iool0, the chans file for the slow channel.


The pinout is as following

-------------
| 1 2 3 4 5 |   Female / Inside View
\  6 7 8 9  /
 \---------/

1 - 11MHz Signal
2 - 30MHz Signal
3 - 55MHz Signal
4 - NC
5 - +5V supply
6 - 11MHz Return
7 - 30MHz Return
8 - 55MHz Return
9 - Supply ground

  5744   Wed Oct 26 23:03:03 2011 kiwamuUpdateLSCRF distribution box : two more 55MHz available

The RF distribution box has been modified so that it generates two more 55 MHz LO signals.

After the modification I put the box back in place.

Then I checked the MICH and YARM locking quickly as a working test of the distribution box and it is working fine so far.

I will update the diagram of the RF distribution box (#4342) tomorrow.

 

(Motivation)

  Since we newly installed POP55 (#5743) an LO signal was needed for the demodulation.

However the RF distribution box didn't have any extra LO outputs.

Therefore we had to make a modification on the RF distribution box so that we can have a 55MHz LO signal for POP55.

Eventually I made two more 55MHz outputs including one spare.

 

(Modification)

The box actually had two extra output SMAs which had been just feed-thru connectors on the front panel without any signals going through.

In the box the modules consist of two categories; the 11MHz system and 55MHz system. I modified only a part of the 55MHz system.

The modification was done in this way:

  * split two branches of 55MHz into four branches by installing two new power splitters (ZMSC-2-1).

  * made and installed some SMA cables whose length were adjusted to be nicely fit in the box.

  * readjustment of the RF levels to 0-2 dBm at the outputs by replacing some attenuators.

  * checked the signals if all of them were happily coming out or not.

Also I found that the POX11 and POY11 demod boards were connected to the whitening filters in a wrong way.

The I and Q signals were in a wrong order. So I corrected them so that the upper inputs in the whitening filter is always the I signal.

 

(RF levels on 55MHz LO outputs)

Since the demod board requires a certain level of the RF signal as as LO, the LO signals have to be 0-2 dBm.

Here are the RF level in each 55MHz output after the adjustment of the level.

     AS55 =  0.76 dBm

     REFL55 = 0.76 dBm

     POP55 = 0.79 dBm

     spare = 0.83 dBm

Those numbers were measured by an oscilloscope and the oscilloscope was configured to measure the rms with the input impedance matched to 50Ohm.

In the measurement I used the actual input seed 55MHz signal from the RF generation box to drive the distribution box.

  5743   Wed Oct 26 20:30:47 2011 KatrinUpdateLSCPOX11 and POP55 installed

[Katrin,Jenne]

RF photo diodes POP55 and POX11 are installed. The beams are aligned to the photo diodes.

 

PD DC out dark DC out bright light power calculated DC output  
POX11 0.1mV 1.3mV 0.09mW 3mV
POP55 35mV 55mV 3 to 4 µW 25mV

I used 0.7 A/W for the response and 50V/A for POP55 according to elog page #4576.

 

To install the third RF photo diode we need to order a plano-convex lens with a focal length of 750 or

maybe even better 1000

  5742   Wed Oct 26 11:35:08 2011 KatrinUpdateGreen LockingYARM PDH box

PDH_w_wo_jump.png

From time to time the 20 dB jump in the transfer function still occurs. The new AD8336 op amp did not change that issue. I am sure that the op amp was broken,

because the amplitude of the sine did not change when I turned the gain knob.

The above two curves were measured with different input amplitude of the sine from the spectrum analyzer. Nothing changed in between except that there was no

jump when Kiwamu was around. Very strange. Testing the electronic board led to no clue what is happening.

For now, I will just use the PDH box as it is, but one should keep this odd behaviour in mind.

  5741   Wed Oct 26 10:07:16 2011 steveUpdateGeneralno electricity ....... will be recheduled

Quote:
Subject: Electricity Interruption- 11/19/11
From: "PP Service Center" <PPService@caltech.edu>
To: "DeLaRosa, Dario" <Dario.DeLaRosa@caltech.edu>,
         ...snip... "Sullivan, Joan O." <sully@caltech.edu>

 

CALIFORNIA INSTITUTE OF TECHNOLOGY

                   FACILITIES MANAGEMENT

              UTILITY & SERVICE INTERRUPTION

 

**PLEASE POST**

 

 

Building:            Central Engineering Services (C.E.S.)

LIGO Gravitational Physics building adjacent to C.E.S.

Safety Storage adjacent to CES

Steele House 

Keck Lab

 

Date:                November 19, 2011

 

Time:                8:00 a.m. To 9:00 a.m.

 

Interruption:        Electricity

 

Contact:             Mike Anchondo ext. 4999  Tom Brennan 4984

 

*This interruption is required for maintenance of high voltage switchgear in Campus Sub Station.

 

 

CANCELLED, WILL BE RESCHEDULED*

  5740   Tue Oct 25 21:49:13 2011 DenUpdateAdaptive FilteringAdaptive filter witness and EP SNR

Quote

Coherence of seismometers to MCL:


STS1 is located at the vertex. x-axis along the x arm.
GUR1 is located at the IMC MC2 mirror. Same orientation.
Coherence.png

=> 1. Only the x-direction has good coherence (to be expected)
     2. Only good coherence at 1.5-4Hz (huh?)

So probably other noise sources are dominating. Let's look into noise projections. Remember IMC autoalignment is off.

A quick adaptive filter run with only the GUR1 and STS1 witnesses applied only to MCL didn't really do anything. Some more thought needs to be invested into the AA and shaping filters.

Indeed, only GUR1_X is reasonable. Static Wiener filtering (length = 2500) of MCL with witness channels GUR_1_X, GUR_1_Y, GUR_1_Z proves your measurements.

We need to callibrate seimometers. I think that now we see velocity, not displacement. It might be useful to amplify the seimometer singal before ADC to make sure that our signal is not ADC noise.

Attachment 1: gur1_x.jpg
gur1_x.jpg
Attachment 2: gur1_y.jpg
gur1_y.jpg
Attachment 3: gur1_z.jpg
gur1_z.jpg
  5739   Tue Oct 25 21:23:17 2011 DmassBureaucracyelogElog Restarted

Elog went nonreponsive. SSH'ed into nodus to run restart script. Elog came back ~15 minutes later.

  5738   Tue Oct 25 20:04:40 2011 MirkoUpdateAdaptive FilteringAdaptive filter witness and EP SNR

We currently have the code running for all DOFs using all witness channels. By default nothing is applied. C-Code parameters can be changed via the respective EPICS variables. Sanity checks in the C-Code make sure the code doesn't crash when nothing / zeros are fed to the code. Let's look into applying FF to one DOF only as a starting point. We start with MCL.

Remember there are two possible signals to look into MC-F and MC-Servo. See page 5695 http://nodus.ligo.caltech.edu:8080/40m/5695

Dark noise: MC-F over MC-Servo which is unconnected in this measurement:
MC-F_SNR_to_Dark_noise.png

=> At least 20dB SNR. ADC noise should not be an issue. Of course more is always better.

Coherence of seismometers to MCL:


STS1 is located at the vertex. x-axis along the x arm.
GUR1 is located at the IMC MC2 mirror. Same orientation.
Coherence.png

=> 1. Only the x-direction has good coherence (to be expected)
     2. Only good coherence at 1.5-4Hz (huh?)

So probably other noise sources are dominating. Let's look into noise projections. Remember IMC autoalignment is off.

A quick adaptive filter run with only the GUR1 and STS1 witnesses applied only to MCL didn't really do anything. Some more thought needs to be invested into the AA and shaping filters.

  5737   Tue Oct 25 18:50:22 2011 SureshUpdateIOOC1IOO model modified to include new WFS lockin structure

Some small fixes to the c1ioo model.

1) I edited the WFS lockin modules to make use of new library part called demod.

2) c1ioo model has been compiled and restarted.

3) fb was restarted at Tue Oct 25 18:43:55 PDT 2011

 

Quote:

I forgot to mention another change I made to the C1IOO model.

The location of the WFS global switch has been shifted. It now cuts off signals just before the WFS servo filters.

I have also added some test points just before the switch so that we can monitor the WFS sensor signals even if the switch is off.

 

 

Quote:

A while back we faced the problem that when we use several lockins to excite the MC degrees of freedom, their relative phase was not known.  The solution suggested was to use one oscillator and several demodulators.

I have now modified the C1IOO.mdl so that this can be implemented.  Previously we were using the MC_ASS lockins for WFS work.  I have not separated the WFS and MC_ASS structures. 

Other jobs to be done in this context are:

1) The medm screens associated with WFS lockins need to be updated with new channel names.

2) The scripts associated with both MC_ASS decentering and WFS ouput matrix determination have to be updated with the new channel names.

3) I also deleted all mdem screens in the $screens$/c1ioo/ directory after copying them to $screens$/c1ioo/bak/.  After installing the new c1ioo model $screens$/c1ioo model now contains just the automatically created screens.  All other user made screens should to into $screens$/c1ioo/master/ directory

This is a pic of the new c1ioo model:

 

 c1ioo20111025.png

 

 

 

  5736   Tue Oct 25 18:09:44 2011 jamieUpdateCDSNew DEMOD part

I forgot to elog (bad Jamie) that I broke out the demodulator from the LOCKIN module to make a new DEMOD part:

DEMOD.png

The LOCKIN part now references this part, and the demodulator can now be used independently.  The 'LO SIN in' and 'LO COS in' should receive their input from the SIN and COS outputs of the OSCILLATOR part.

  5735   Tue Oct 25 16:24:58 2011 SureshUpdateIOOC1IOO model modified to include new WFS lockin structure

I forgot to mention another change I made to the C1IOO model.

The location of the WFS global switch and the WFS_GAIN have been shifted. The switch now cuts off signals just before the WFS servo filters.

I have also added some test points just before the switch and the so that we can monitor the WFS error signals which would be unaffected even if the WFS_GAIN is changed..

 

 

Quote:

A while back we faced the problem that when we use several lockins to excite the MC degrees of freedom, their relative phase was not known.  The solution suggested was to use one oscillator and several demodulators.

I have now modified the C1IOO.mdl so that this can be implemented.  Previously we were using the MC_ASS lockins for WFS work.  I have not separated the WFS and MC_ASS structures. 

Other jobs to be done in this context are:

1) The medm screens associated with WFS lockins need to be updated with new channel names.

2) The scripts associated with both MC_ASS decentering and WFS ouput matrix determination have to be updated with the new channel names.

3) I also deleted all mdem screens in the $screens$/c1ioo/ directory after copying them to $screens$/c1ioo/bak/.  After installing the new c1ioo model $screens$/c1ioo model now contains just the automatically created screens.  All other user made screens should to into $screens$/c1ioo/master/ directory

This is a pic of the new c1ioo model:

 

 c1ioo20111025.png

 

 

  5734   Tue Oct 25 11:48:02 2011 KatrinHowToElectronicssolder tiny smd op amps

Yesterday, I had the great pleasure to solder a tiny 4 x 4 mm op amp with 16 legs (AD8336).

I figured out that the best and fastest way to do it is

  1. to put solder with the soldering iron on every contact of the electronic board (top side)
  2. heat the bottom side of the electronic board with a heat gun
  3. use a needle to test if the solder is melted
  4. if it is melted place the op amp on the electronic board
  5. apply some vertical force on the op amp for proper contact and heat for 1 to 2 more minutes
  6. done

 

  5733   Tue Oct 25 01:19:17 2011 SureshUpdateComputersfb restarted and c1ioo model committed to svn

When I installed the new model I restarted the fb between 1 and 1:30 AM PDT Oct 25, 2011

  5732   Tue Oct 25 01:14:15 2011 SureshUpdateIOOC1IOO model modified to include new WFS lockin structure

A while back we faced the problem that when we use several lockins to excite the MC degrees of freedom, their relative phase was not known.  The solution suggested was to use one oscillator and several demodulators.

I have now modified the C1IOO.mdl so that this can be implemented.  Previously we were using the MC_ASS lockins for WFS work.  I have now separated the WFS and MC_ASS structures. 

Other jobs to be done in this context are:

1) The medm screens associated with WFS lockins need to be updated with new channel names.

2) The scripts associated with both MC_ASS decentering and WFS ouput matrix determination have to be updated with the new channel names.

3) I also deleted all medm screens in the $screens$/c1ioo/ directory after copying them to $screens$/c1ioo/bak/.  After installing the new c1ioo model $screens$/c1ioo directory now contains just the automatically created screens.  All other user made screens should go into $screens$/c1ioo/master/ directory

This is a pic of the new c1ioo model:

 

 c1ioo20111025.png

 

  5731   Mon Oct 24 20:00:21 2011 MirkoUpdateCDSTiny little scripts

Located in /opt/rtcds/caltech/c1/userapps/release/cds/common/scripts

Script 1: Diagreset.sh
Hits the diag reset buttons on the models on c1lsc and c1sus computers.

Script 2: Burtrest.sh
Restores the burt files from "today" 4am. Use a text editor if you want to change the times.

  5730   Mon Oct 24 19:48:16 2011 MirkoUpdateAdaptive FilteringFilter execution time

Toyed around some more with the adaptive filters.

Execution time:

nTaps    Downsampling factor     Execution time average / max in ca. 3 min [us], (480 us available)
1000     16                                110 / 150
2000     16                                280 / 340
3000     16                                380 / 470
4000     16                                Over limit

Now we are running with Downsampling 32, 4000 Taps => max 410us execution time.

I tried to desynchronize the downsampled operations of the filters of the different DOFs. That however increased execution time by about 10%. So I undid that.

 

  5729   Mon Oct 24 17:23:14 2011 SUS_DiagonalizerUpdateSUSOptics kicked
This is a cron-elog test. No optics have been kicked.
  5728   Mon Oct 24 14:26:01 2011 steveUpdateGeneralno electricity on Nov. 19
Subject: Electricity Interruption- 11/19/11
From: "PP Service Center" <PPService@caltech.edu>
To: "DeLaRosa, Dario" <Dario.DeLaRosa@caltech.edu>,
         ...snip... "Sullivan, Joan O." <sully@caltech.edu>

 

CALIFORNIA INSTITUTE OF TECHNOLOGY

                   FACILITIES MANAGEMENT

              UTILITY & SERVICE INTERRUPTION

 

**PLEASE POST**

 

 

Building:            Central Engineering Services (C.E.S.)

LIGO Gravitational Physics building adjacent to C.E.S.

Safety Storage adjacent to CES

Steele House 

Keck Lab

 

Date:                November 19, 2011

 

Time:                8:00 a.m. To 9:00 a.m.

 

Interruption:        Electricity

 

Contact:             Mike Anchondo ext. 4999  Tom Brennan 4984

 

*This interruption is required for maintenance of high voltage switchgear in Campus Sub Station.

 

 

  5727   Fri Oct 21 18:20:54 2011 MirkoUpdateCDSFirst OAF version running

[Jenne, Jamie, Mirko]

We got the first version of the oaf code based on Matt"s code running!! :-)
Produces already data for e.g. MICH DOF. But don"t trust that. It's only 10 taps long and delay is not adjusted.

  5726   Fri Oct 21 16:59:14 2011 KatrinUpdateGreen LockingYARM PDH box broken

I could not improve the locking. So, I checked the transfer function of the PDH box again. The transfer function looks okay if the gain knob is <=2.0.

If the gain knob is >2.0 the 20dB step appears in the transfer function (see elog page 5713). This step is shifted to higher frequencies if the gain is

increased. The PZT drive out was not saturated at any time. Yesterday, I checked the electronic circuit with a gain of 2.0. Thus,  I couldn't find the broken

gain amplifier (AD8336). The amplifier is ordered in will arrive on Monday.

  5725   Fri Oct 21 16:06:12 2011 SureshUpdateComputer Scripts / ProgramsMC input matrices empty again

The MC suspensions were not damping and the reason was traced to the empty imput matrices in the suspension controls.  This has been an issue in the past as well when the sus machine is rebooted some of the burt restore does not populate these matrices.

I ran the burtgooey and restored c1mcs.snap file.  from a couple of hours ago.

 

 

  5724   Fri Oct 21 15:49:35 2011 SureshUpdateIOOPMC input alignment improved

The image on the PMCR camera was quite assymetric and PMC output was at 80% .... upon improving the alignment I managed to push it up to 87%

 

  5721   Fri Oct 21 11:02:47 2011 steveUpdatePSLPSL laser turned ON

Quote:

In order to move the emergency shut off switch in room 103 I had to turn off the 2 W Innolight laser. This job will take an hour.

 

 It is back on.

Attachment 1: laserofon.png
laserofon.png
  5720   Fri Oct 21 10:32:14 2011 steveUpdatePSLPSL laser turned off

In order to move the emergency shut off switch in room 103 I had to turn off the 2 W Innolight laser. This job will take an hour.

 

  5719   Fri Oct 21 09:08:33 2011 steveUpdateSUSMC2 Misalignment

 Thinks to do before the NEXT realignment:

B,  tie 4 ancher bolts on table legs to the floor

C,  tie 4 dog clamps between table and chamber

D,  check the locked position of the 4 x 4 positioning screws

E,  check bellow protecting 4 tubes are not shorting

A,  here is the concrete  slab cut

 

It reminds me to check the  IFO vacuum envelope dog clamps on the chambers to floor  with torque wrench.

 

 

Attachment 1: mc2.jpg
mc2.jpg
  5718   Fri Oct 21 02:57:38 2011 SureshSummarySUSMC2 Misaligned 2:27PM on Wednesday : cause traced

Quote:

Quote:

There looks some activity at around MC2 on Wednesday afternoon.
It caused the misalignment of MC2. Misalignment was not found in MC1/3.

It seems that the incident beam on the MC was aligned in the evening.
This increased the MC transmission but it is vibible that the spot on MC2 is shifted from the center.

We need an action on this issue tomorrow in the daytime.

 

I am working on fixing this.  You might some strange stuff going on in the control room screens.  Pls ignore it till I am done.

 

While chatting with Jenne I learnt that some substantial amount of work had taken place yesterday around the MC2 chamber.  This was associated with the relocating of seismometers.  ref elog

I reiterate what is well known for quite sometime:  MC2 table is not well isolated from the ground.  And we should not approach this chamber unless absolutely necessary. I have blocked off the area around it which we should avoid.  It is a serious waste of time and effort to realign the MC each time the MC2 table decides to settle into a new position.

Steve tells me that the mild-steel frame supporting the chamber+MC2_table sits with two legs on one concrete slab while the other two legs sit on another one.   The frame is also quite weak without sufficient gussets or cross connects.  The next time we have a major shutdown we must replace this frame with a more sturdy one which sits on one slab (preferably the one on which the rest of the MC sits).

Till we improve this mounting, I suggest that we avoid that area as much as possible.

 

  5717   Fri Oct 21 02:36:44 2011 SureshSummarySUSMC2 Misaligned 2:27PM on Wednesday : MC Realigned

Quote:

Quote:

There looks some activity at around MC2 on Wednesday afternoon.
It caused the misalignment of MC2. Misalignment was not found in MC1/3.

It seems that the incident beam on the MC was aligned in the evening.
This increased the MC transmission but it is vibible that the spot on MC2 is shifted from the center.

We need an action on this issue tomorrow in the daytime.

 

I am working on fixing this.  You might some strange stuff going on in the control room screens.  Pls ignore it till I am done.

 

 

  I have realigned the MC by recentering the spots on all the MC optics.  The current spot positions (in mm) are:

MC1P     MC2P     MC3P      MC1Y      MC2Y     MC3Y

0.2245    0.3364   -0.2801   -1.8891    0.1631   -1.744

Initially the lockins 2 and 5 showed very small outputs.  This was traced to the fact that we have recently switched on a 28Hz ELP filter module in the MC2 ASC filter bank which introduces an extra phase of about 75deg..  See this elog.

When the MC ASS lockins were initially setup, the phase was set with this filter module switched off.  Since quite some time has passed since the last calibration of these phases, I readjusted the phases to minimise the  Q_OUTPUT and I also adjusted the GAINs in the SIG filter banks  of all the six lockins so that their I_OUT's drop by the calibration value of -2.65 when an offset of 0.1 is introduced into the MC suspension output matrices.  Two short scripts in the $scripts$/ASS/ directory help in setting and removing these offsets.  They are called MCxoffsetOn and MCxoffsetOff.   They have to be edited appropriately to address each DoF of the MC.

The $scripts$/ASS/mcassUp script., which sets up everything to make the MC spot decentering measurement, has been edited to set these new phases and gains.  The old settings have been commented out.

I then centered the spots on the WFS sensors and the MC_TRANS QPD.  We are now ready to make the MC WFS output matrix transfer coef measurement again, but this time with the WFS loops closed.

 

  5716   Thu Oct 20 18:57:35 2011 SureshSummarySUSMC2 Misaligned 2:27PM on Wednesday

Quote:

There looks some activity at around MC2 on Wednesday afternoon.
It caused the misalignment of MC2. Misalignment was not found in MC1/3.

It seems that the incident beam on the MC was aligned in the evening.
This increased the MC transmission but it is vibible that the spot on MC2 is shifted from the center.

We need an action on this issue tomorrow in the daytime.

 

I am working on fixing this.  You might some strange stuff going on in the control room screens.  Pls ignore it till I am done.

 

  5715   Thu Oct 20 18:42:47 2011 KojiUpdateComputer Scripts / ProgramsWhere should the "Update Snapshots" of screens live?

The following directory exists. We can apply this convention to all of the models.

/cvs/cds/rtcds/caltech/c1/medm/c1lsc/snap

Quote:

While trying to implement the regular yellow shell script button in MEDM for my new OAF screen, I noticed that the update snapshot stuff in all of the buttons that I checked (including IFO Align and LSC Overview) are pointing to folders in the old /cvs/cds/caltech/ area.  Also, I think some of the folders that it's looking for don't exist anymore, even in the old system.  So.  Has anyone thought about where the snapshots should live in the new world order?  Previously they were in ...../medm/c1/subsystem/ .  Maybe we should make a snapshots folder in each subsystem's medm folder, at the same level as the 'master' folder for the custom screens?  This is my current proposal.

Unless someone objects / has a better plan / knows why they're still pointing to the old place, I'll do this in the morning, and work on changing all the buttons to point to the new place.

 

  5714   Thu Oct 20 18:01:17 2011 JenneUpdateComputer Scripts / ProgramsWhere should the "Update Snapshots" of screens live?

While trying to implement the regular yellow shell script button in MEDM for my new OAF screen, I noticed that the update snapshot stuff in all of the buttons that I checked (including IFO Align and LSC Overview) are pointing to folders in the old /cvs/cds/caltech/ area.  Also, I think some of the folders that it's looking for don't exist anymore, even in the old system.  So.  Has anyone thought about where the snapshots should live in the new world order?  Previously they were in ...../medm/c1/subsystem/ .  Maybe we should make a snapshots folder in each subsystem's medm folder, at the same level as the 'master' folder for the custom screens?  This is my current proposal.

Unless someone objects / has a better plan / knows why they're still pointing to the old place, I'll do this in the morning, and work on changing all the buttons to point to the new place.

  5713   Thu Oct 20 16:33:24 2011 KatrinUpdateGreen LockingTransfer function YARM PDH box

Yesterday, I measured the transfer function of the YARM PDH box.

SCRN0000.pdf

 

I tested the electronic board and couldn't find a frequency dependent behaviour. So I measured the TF again and it looked nice.

PDH_box.png

Today's nice measurement could is/was reproducible. I suppose yesterday's measurement is just an artefact.

The electronic board is modified according to Kiwamu's wiki entry http://blue.ligo-wa.caltech.edu:8000/40m/Electronics/PDH_Universal_Box

 

Btw. The light could be locked to the cavity for ~3min.

  5712   Thu Oct 20 12:43:19 2011 steveUpdateSAFETYrefilled first aid kits & their locations

First aid kits are located close vicinity of entry doors and under circuit breaker panels.

 

Attachment 1: P1080299.JPG
P1080299.JPG
Attachment 2: P1080306.JPG
P1080306.JPG
Attachment 3: P1080304.JPG
P1080304.JPG
Attachment 4: P1080305.JPG
P1080305.JPG
Attachment 5: P1080300.JPG
P1080300.JPG
Attachment 6: P1080301.JPG
P1080301.JPG
Attachment 7: P1080303.JPG
P1080303.JPG
Attachment 8: P1080307.JPG
P1080307.JPG
  5711   Thu Oct 20 11:59:21 2011 ZachUpdateComputer Scripts / Programsmodified "dataviewer" on nodus

The "dataviewer" script was still setting the server to fb40m on nodus. I modified it to fb, so that this is the default when you enter "dataviewer" or "dv".

  5710   Thu Oct 20 09:54:53 2011 jamieUpdateComputer Scripts / Programspynds working on pianosa again

Quote:

 

Doesn't work on pianosa either. Has someone changed the python environment?

pianosa:SUS_SUMMARY 0> ./setSensors.py 1000123215 600 0.1 0.25
Traceback (most recent call last):
  File "./setSensors.py", line 2, in <module>
    import nds
ImportError: No module named nds

 So I found that the NDS2 lib directory in (/ligo/apps/nds2/lib) was completely empty.  I reinstalled NDS2 and pynds, and they are now available again by default on pianosa (it should "just work", assuming you don't break your environment).

Why the NDS2 lib directory was completely empty is definitely a concern to me.  The contents of directories don't just disappear.  I can't imagine how this would happen other than someone doing it, either on purpose or accidentally.  If someone actually deleted the contents of this directory on purpose they need to speak up, explain why they did this, and come see me for a beating.

  5709   Thu Oct 20 04:47:37 2011 kiwamuUpdateLSCclipping search round 1

[Koji / Kiwamu]

  We tried finding a possible clipping in the vertex part.

We couldn't find an obvious location of a clipping but found that the recycling gain depended on the horizontal translation of the input beam.

We need more quantitative examination and should be able to find a sweet spot, where the recycling gain is maximized.

 

(what we did)

  + locked the carrier-resonant PRMI.

  + with IR viewers we looked at the inside of ITMX, ITMY and BS chambers to find an obvious clipping.

    => found two suspicious bright places and both were in the ITMY chamber.

      (1) POY pick off mirror : looked like a small portion of a beam was horizontally clipped by the mirror mount but not 100% sure whether if it is the main beam or a stray beam.

      (2) The top of an OSEM cable connectors tower : although this is in the way of the SRC path and nothing to do with PRC.

 + Made a hypothesis that the POY mirror is clipping the main beam.

 + To reject/prove the hypothesis we shifted the translation of the incident beam horizontally such that more beam hits on the suspicious mirror

 + Realigned and relocked PRMI.

    => Indeed the recycling gain went down from 6 to 0.8 or so. This number roughly corresponds to a loss of about 50%.

         However the MICH fringe still showed a very nice contrast (i.e. the dark fringe was still very dark).

         Therefore our conclusion is that the POY mirror is most likely innocent.

  5708   Thu Oct 20 01:40:33 2011 KojiSummarySUSMC2 Misaligned 2:27PM on Wednesday

There looks some activity at around MC2 on Wednesday afternoon.
It caused the misalignment of MC2. Misalignment was not found in MC1/3.

It seems that the incident beam on the MC was aligned in the evening.
This increased the MC transmission but it is vibible that the spot on MC2 is shifted from the center.

We need an action on this issue tomorrow in the daytime.

Attachment 1: MC2_misalign.png
MC2_misalign.png
  5707   Wed Oct 19 19:43:16 2011 JenneUpdatePSLPMC found unlocked

Quote:

I just relocked the PMC.  I don't know why it was unlocked.

 Again....

  5706   Wed Oct 19 18:18:03 2011 SureshUpdateElectronicsStochMon : Filters installed

Quote:

To get to the bottom of the RFAM mystery, we've got to resurrect the StochMon to trend the RFAM after the IMC.

We will put an 1811 on the MC_TRANS or IP_POS beam (the 1811 has an input noise of 2.5 pW/rHz).

Then the Stochmon has an input pre-amp, some crappy filters, and then Wenzel RMS->DC converters. We will replace the hand-made filters with the following ones from Mini-Circuits which happen to match our modulation frequencies perfectly:

11 MHz     SBP-10.7+

55 MHz     SBP-60+

29.5 MHz   SBP-30+

 

The Stochmon had a four-way splitter, four hand-made filters and four mini-circuits ZX47-60-S+ Power Detectors.

Using the filters from our stock I have replaced the hand-made filters with the ones mentioned in Rana's elog.   The power supply solders to the ZX47-60-S+ Power Detectors were weak and came off during reassembly. And some of the handmade short SMA cables broke off at the neck.  So I changed the power supply cables and replaced all the short SMA cables with elbows.  I also removed one of the Power Detectors since there were four in the box and we need only three now.

The power supply connector on the box is illegal.  The current lab standard for  {+15, 0 , -15}  uses that connector.  So we are going to change it as soon as possible.  We need to identify a good {0, 5} lab standard and stock them.

The following were removed from the box:

PA190146.JPG

 

The box now looks like this:

PA190143.JPG

 

Steps remaining in installation of Stochmon:

1) Install the Newfocus 1811 PD at the IPPOS by diverting some of the power in that path

2) Connect the outputs of the Stochmon to ADC inputs in 1X2 rack.

ELOG V3.1.3-